# Class-Incremental Learning using Diffusion Model for Distillation and Replay

### v1

@@ 1 Introduction
In a class-incremental learning (CIL) scenario, deep neural network models are not trained offline on a fixed pre-collected dataset but sequentially on new incoming data. The model has to be incrementally updated using only a limited number of new classes at a time, with the objective of learning a unified classifier among all seen classes. This paradigm, similar to the human learning process, is challenging as models are suffering from catastrophic forgettingĀ [19, 21]: learning new data or classes will result in severe degradation of the performance on the past ones.
Multiple researchers highlighted the importance of rehearsal for continual and incremental learning: by storing samples from each previously encountered class in a memory, models are able to replay past data while learning new ones in order to mitigate catastrophic forgetting. However, the size of this memory is often limited to a few samples per class due to either storage limitation or privacy concerns. Several authorsĀ [14, 15, 17, 33] have proposed to leverage additional external training data as a mean to work around this constraint. Those additional training data are sampled from large curated datasets of real images, such as ImageNet, and belong to classes different from the ones encountered by the incremental learner. They are mostly used in an unsupervised manner, for distillating knowledge from the past model to the new one.
In this work, we propose to use the state-of-the-art generative model Stable DiffusionĀ [23] as a source of complementary data for class-incremental learning. Compared to approaches that rely on additional external real datasets, our method generates synthetic samples belonging to the same classes as the ones previously encountered by the model. This fundamental difference allows us to make use of those samples not only for the knowledge distillation loss but also for replay in the classification loss whereas concurrent works only focus on the distillation loss. Ablation study highlights the improvement resulting from using the additional data for both losses. To the best of our knowledge, this is the first attempt on using large pretrained text-to-image generative models to improve general class-incremental learning methods on large scale datasets.
@@ 2 Related work
Class-incremental Learning (CIL): Methods for class-incremental learningĀ [34] usually rely on knowledge distillation and a small replay memory containing past samples combined with a bias mitigation method. Knowledge distillation is used to preserve previous knowledge, it may be logit distillation Ā [16, 22, 31], feature distillationĀ [8] or relational distillationĀ [28]. Using a memory for replaying previously learned exemplars is a simple yet effective method to recall old classes knowledge. However, the limited size of the memory induces a bias toward the new classes that strongly degrade the performance of the model if not mitigated. To this end, several methods have been proposed such as Nearest-Mean-of-Exemplars (NME) classifierĀ [22], cosine classifierĀ [8], bias correction layerĀ [31], finetuning on a balanced subsetĀ [3] or specific lossesĀ [1, 11, 12]. To overcome the limitation induced by the limited size of the memory, several works have also proposed to use model inversionĀ [27, 32] or to train a generative model in parallel to generate samples from past classesĀ [4, 26, 30]. These methods rely on the quality of the generated data which is challenging for large scale datasets and are impacted by catastrophic forgetting when trained sequentially.
@@
Additional data for CIL: Various approaches have studied the use of additional external data, either as a direct substitute in case of the absence of memoryĀ [15, 33] or as a complementary resource during trainingĀ [14, 17]. Global distillation (GD)Ā [15] proposed a confidence-based sampling strategy to sample additional external data and used them for a specifically-designed distillation loss. Deep Model Consolidation (DMC)Ā [33] consists in training one distinct model for each training step and use external data to combine them into a unified model. LechatĀ et al.Ā [14] proposed a pseudo-labeling approach to better exploit the additional external unlabeled data for representation learning. LiuĀ et al.Ā [17] observed that using samples from new classes in the knowledge distillation loss negatively impacts the training of incremental models and proposed to replace them using placebos of old classes sampled from an external dataset using an online learning method. Those different approaches rely solely on real images, most of the time unlabeled, by sampling their additional data from large curated datasets: primarily ImageNetĀ [24] or TinyImages datasetĀ [29].
@@
Learning from synthetic data: In one of the pioneer works on learning from synthetic images, BesnierĀ et al.Ā [2] trained a classifier for 10 classes of ImageNet using samples generated by a class-conditional GAN trained on ImageNet. JahanianĀ et al.Ā [10] used a GAN to generate multiple views for contrastive learning methods. Recently, following the recent breakthrough in text-to-image generative models, SariyildizĀ et al.Ā [25] proposed the use of those general-purpose text-conditioned pretrained generative models for synthesizing datasets as a direct replacement of real image datasets for the training of large scale image-level classification models. While these models can not reach the same accuracy than the models trained using real images when tested on real test datasets, they exhibit other advantages such as a more transferable representation. Concurrently, HeĀ et al.Ā [6] also proposed to use large synthetic datasets for improving pretrained models on zero-shot and few-shot learning, and transfer learning.
@@
In this work, we propose to study the use of widely available pretrained generative models such as Stable DiffusionĀ [23] as a source of additional labeled samples combinable with general class-incremental learning methods applied on large-scale datasets.
@@ 3 Proposed method
The training procedure is divided into steps: the base step, often referred to as step or initial step, followed by incremental steps. Each step consists of a training dataset containing images belonging to the set of new and previously unseen classes ; for . During each step, the model is trained on where is the replay memory containing few samples from previously encountered classes. After each step, the model is evaluated on the test set of all the classes learned so far without having access to any step or task descriptor. Additionally, during every step, the model also has access to an external source of data that may or may not change during the training. This external source of additional data can either be an online, and potentially infinite stream of data that the model can fetch when needed or a fixed set of data accessible in an offline manner.
@@ 3.1 Synthetic images
Following the recent in-depth study of SariyildizĀ et al.Ā [25] we propose to use a pretrained Stable DiffusionĀ [23] as our source of additional data for class-incremental learning.
Prompting: To generate a sample for , the generative model Stable Diffusion requires a textual prompt. Similarly toĀ [25], to generate a synthetic image for the class c we use the prompt āā where āā is the textual name of the class and āā its description. Although using only ācā as a prompt may be enough in most cases, this may lead to semantic errors with homographs. For example for CIFAR100 dataset, using only āappleā as a prompt may generate images of products or shops from the homonym brand instead of the fruit. Adding the definition of the class to the prompt helps to solve this issue. To produce these prompts automatically without any further engineering, WordNetĀ [20] is used to provide the lemmas of the synset as the class name ācā and the definition of the synset as the description āā. For datasets such as ImageNet, the association of each class with a synset of WordNet is included in the dataset itself. For other datasets such as CIFAR100, the association with WordNet can be done semi-automatically at a negligible cost. FigureĀ 1 shows synthetic samples for CIFAR100 before being resized to 32x32. In the cases where the mapping between the target dataset and WordNet can not be done, it may be necessary to either manually design and tune the prompt for each class or to rely on a different automation approach such as using image-to-text model for example.
Guidance scale: The classifier-free guidanceĀ [7] controls the trade-off between the quality of the generated synthetic samples and their diversity in diffusion models. Following SariyildizĀ et al.Ā [25], we also used a guidance scale of compared to the default to increase the diversity of our source of additional data at the cost of a slight decrease of the quality of each sample.
@@ 3.2 SDDR for Class-incremental learning
| iCaRLĀ [22] | 57.17 | 52.57 | - | 65.04 | 59.53 | - | 51.36 | 46.72 |
| BiCĀ [31] | 59.36 | 54.20 | 50.00 | 70.07 | 64.96 | 57.73 | 62.65 | 58.72 |
| LUCIRĀ [8] | 63.42 | 60.18 | - | 70.47 | 68.09 | - | 64.34 | 61.28 |
| LUCIR w/ MnemonicsĀ [18] | 63.34 | 62.28 | 60.96 | 72.58 | 71.37 | 69.74 | 64.54 | 63.01 |
| LUCIR w/ DDEĀ [9] | 65.27 | 62.36 | - | 72.34 | 70.20 | - | 67.51 | 65.77 |
| LUCIR w/ BalancedS-CEĀ [12] | 64.83 | 62.36 | 58.38 | 71.18 | 70.66 | 65.10 | 67.81 | 66.47 |
| TPCILĀ [28] | 65.34 | 63.58 | - | 76.27 | 74.81 | - | 64.89 | 62.88 |
@@
Using the previously described generative model, we propose a new method for class-incremental learning: Stable Diffusion for Distillation and Replay (SDDR).
@@
Datasets of real images used by concurrent approaches as an external source of additional data for class-incremental learning do not contain any image belonging to the same classes as the one encountered by the incremental model. This leads these methods to mainly use this external source of data in an unsupervised manner for knowledge distillation only. By leveraging the pretrained text-to-image generative model, our method is able to generate an additional dataset containing labeled images belonging to the same classes as the one previously learned by the model without needing a costly manual labeling and curating process. This allows us to use the additional dataset for distillation but also for replay in the classification loss. The ablation study in SectionĀ 4.3 shows the advantages of our approach.
@@
Our approach is designed as a complementary method that can be seamlessly combined with other standard methods for class-incremental learning. AlgorithmĀ 1 describes a training procedure of our approach SDDR when combined with a class-incremental learning model. Each step of the class-incremental learning procedure is divided into two distinct phases. During the first phase, the model is trained following the method it is combined with, using a mixed batch of data. As illustrated in FigureĀ 2, each batch used to train the model is composed of half of synthetic images sampled from and the other half of real data sampled from the memory and the current dataset . Then in the second phase, before moving to the next incremental step, the synthetic dataset is extended by using Stable Diffusion to generate synthetic samples for each new class encountered in the current dataset , following the method detailed in SectionĀ 3.1. Therefore, at the end of each incremental step, the additional dataset would contain samples where is the number of classes encountered up to the step , included. In the ablation study in SectionĀ 4.3, we also analyze the use of the synthetic datasets as a straightforward replacement of the replay memory for the more challenging setting of exemplar-free class-incremental learning.
@@
In our work, we considered that the additional dataset of synthetic samples is an offline dataset stored on the device itself. Nonetheless, if the storage is a constraint for the incremental learner, the additional dataset can be used in an online way as an infinite stream of data, similarly toĀ [17], by either using the generative model locally or by relying on a cloud-based model. Compared to concurrent works relying on real datasets as a source of additional data, our method offers more flexibility depending on the constraint of the problem: it can be used with or without limited storage, computational budget, and capacity to communicate with external services.
@@
The quality of the synthetic images, their fidelity to the real target, and the variety of concepts available in the additional dataset are intrinsically limited by the pretrained generative text-to-image model used. Using a different and more specialized generative model or finetuning it to the specific target training dataset may further improve the performance of the incremental learning model it is combined with.
@@ 4 Experiments
4.1 Experimental setups
Datasets: Experiments are conducted on three datasets: CIFAR100Ā [13], ImageNet-Subset and ImageNetĀ [24]. CIFAR100 is composed of 60,000 32x32 RGB images from 100 classes with 500 training and 100 testing samples per class. ImageNet (ILSVRC 2012) is composed of around 1.28 million high-resolution images from 1,000 classes with about 1,300 training and 50 testing sample per class. ImageNet-Subset is a subset of ImageNet containing only 100 classes. Following the experimental setting initially proposed by HouĀ et al.Ā [8], an initial step containing half of the classes is followed by (5, 10, or 25) incremental steps containing the remaining classes. The class order is defined by NumPy using the random seed 1993. The training images are normalized, randomly horizontally flipped, and cropped, and no more augmentation is applied.
Comparison methods: We combine our proposed method SDDR with two fundamental baseline approaches for class-incremental learning: iCaRLĀ [22] and LUCIRĀ [8]. We selected these two methods because they are representative of the principal approaches for class-incremental learning; using different distillation losses (logits distillation and feature distillation) and difference bias mitigation methods (Nearest-Mean Exemplar classifier and cosine normalized classifier). For comparison, we also include the recent approach PlaceboCILĀ [17] which leverages external unlabeled datasets of real high-resolution images to improve the distillation loss of class-incremental learning methods.
Implementation details: Following the standard settingĀ [8, 22], a 32-layer ResNetĀ [5] is used for CIFAR100 and a 18-layer ResNet for ImageNet and ImageNet-Subset. Every considered method uses a growing memory containing 20 exemplars per class during the training procedure. The number of epochs, weight decay, learning rate, momentum, and learning rate scheduling are identical to HouĀ et al.Ā [8] implementation. During the incremental step, each batch is composed of 128 images sampled from the generative model and 128 images sampled from the union of the new dataset and the replay memory. Some methods for class-incremental learning use other losses in addition to the classification and distillation losses. Using synthetic data in these other losses may improve or degrade the performance of the method on a case-by-case basis. Specifically, in our experiments on LUCIRĀ [8] we decide to not use the synthetic data for the margin ranking loss as preliminary experiments suggested that it slightly decreases the performance of the method. The generative model we use is Stable Diffusion version 1.4Ā [23] with a guidance scale of 2.0, 50 steps and the prompt āāĀ [25]. Synthetic images have been generated using specific seeds before the incremental training and every method are using the same datasets of synthetic images. For CIFAR100, the synthetic dataset contains 500 32x32 images per class, and for ImageNet and ImageNet-Subset 1300 512x512 images per class. For every dataset, synthetic images are generated with size 512x512 and are only resized to 32x32 for CIFAR100 prior to the experiments.
Performance Measure: Models are evaluated and compared using the Average Incremental Accuracy defined by RebuffiĀ et al.Ā [22]. It is the average of the Top-1 accuracy of the model on the test data of all the classes seen so far, at the end of each training step including the initial base step.
4.2 Comparison with baselines
@@
In TableĀ 1, the average incremental accuracy (Top-1) of our method and the different baselines are reported for CIFAR100, ImageNet-Subset, and ImageNet with 5, 10, and 25 incremental steps using a replay memory of 20 real exemplars per class for every method.
For both LUCIR and iCaRL, using our proposed approach SDDR lead to a significant improvement of the average incremental accuracy: from 0.87 percentage points (p.p.) to 6.25p.p. for LUCIR and from 3.3p.p. to 8.44p.p. for iCaRL, depending on the dataset and the setting. This improvement is especially important in more challenging settings with numerous incremental steps or when the replay memory is small as discussed in SectionĀ 4.3.
@@
When looking more closely at the learned model, we can see that our approach improve the capacity of the model to preserve past knowledge without penalizing its plasticity. This is especially true in the most challenging settings with a small replay memory: for example on CIFAR100 with 10 incremental steps and 5 exemplars per class in the memory, by combining LUCIR with SDDR, the final accuracy on the 50 bases classes increases from 34.15% to 42.06% while the accuracy on the 50 remaining classes also increases from 46.08% to 52.23%.

### v2

@@ 1 Introduction
In a class-incremental learning (CIL) scenario, deep neural network models are not trained offline on a fixed pre-collected dataset but sequentially on new incoming data. The model has to be incrementally updated using only a limited number of new classes at a time, with the objective of learning a unified classifier among all seen classes. This paradigm, similar to the human learning process, is challenging as models are suffering from catastrophic forgetting [21, 23]: learning new data or classes will result in severe degradation of the performance on the past ones.
Multiple researches have highlighted the importance of rehearsal for continual and incremental learning: by storing samples from each previously encountered class in a memory, models are able to replay past data while learning new ones in order to mitigate catastrophic forgetting. However, the size of this memory is often limited to a few samples per class due to either storage limitation or privacy concerns. Several authors [16, 17, 19, 37] have proposed to leverage additional external training data as a means to work around this constraint. Those additional training data are sampled from large curated datasets of real images, such as ImageNet, and belong to classes different from the ones encountered by the incremental learner. They are mostly used in an unsupervised manner, for distillating knowledge from the past model to the new one.
In this work, we propose to use the pretrained state-of-the-art generative model Stable Diffusion [25] as a source of complementary data for class-incremental learning. Compared to approaches that rely on additional external real datasets, our method generates synthetic samples belonging to the same classes as the ones previously encountered by the model. This fundamental difference allows us to make use of those samples not only for the knowledge distillation loss but also for replay in the classification loss whereas concurrent works only focus on the distillation loss. Ablation study highlights the improvement resulting from using the additional data for both losses. To the best of our knowledge, this is the first attempt to use large pretrained text-to-image generative models to improve general class-incremental learning methods on large scale datasets.
@@ 2 Related work
Class-incremental Learning (CIL): Methods for class-incremental learning [38] usually rely on knowledge distillation and a small replay memory containing past samples combined with a bias mitigation method. Knowledge distillation is used to preserve previous knowledge, it may be logit distillation [18, 24, 34], feature distillation [5, 10] or relational distillation [30]. Using a memory for replaying previously learned exemplars is a simple yet effective method to recall knowledge of old classes. However, the limited size of the memory induces a bias toward the new classes that strongly degrade the performance of the model if not mitigated. To this end, several methods have been proposed such as Nearest-Mean-of-Exemplars (NME) classifier [24], cosine classifier [10], bias correction layer [34], finetuning on a balanced subset [3] or specific losses [1, 13, 14]. To overcome the limitation induced by the limited size of the memory, several works have also proposed to use model inversion [29, 36] or to train a generative model in parallel to generate samples from past classes [6, 28, 33]. These methods rely on the quality of the generated data, which poses a challenge for large scale datasets and is impacted by catastrophic forgetting when trained sequentially.
@@
Additional data for CIL: Various approaches have studied the use of additional external data, either as a direct substitute in case of the absence of memory [17, 37] or as a complementary resource during training [16, 19]. Global distillation (GD) [17] proposed a confidence-based sampling strategy to sample additional external data and used them for a specifically-designed distillation loss. Deep Model Consolidation (DMC) [37] consists in training one distinct model for each training step and use external data to combine them into a unified model. Lechat et al. [16] proposed a pseudo-labeling approach to better exploit the additional external unlabeled data for representation learning. Liu et al. [19] observed that using samples from new classes in the knowledge distillation loss negatively impacts the training of incremental models and proposed to replace them using placebos of old classes sampled from an external dataset using an online learning method. Those different approaches rely solely on real images, most of the time unlabeled, by sampling their additional data from large curated datasets: primarily ImageNet [26] or TinyImages dataset [31].
@@
Learning from synthetic data: In one of the pioneer works on learning from synthetic images, Besnier et al. [2] trained a classifier for 10 classes of ImageNet using samples generated by a class-conditional GAN trained on ImageNet. Jahanian et al. [12] used a GAN to generate multiple views for contrastive learning methods. Recently, following the recent breakthrough in text-to-image generative models, Sariyildiz et al. [27] proposed the use of those general-purpose text-conditioned pretrained generative models for synthesizing datasets as a direct replacement of real image datasets for the training of large scale image-level classification models. While these models can not reach the same accuracy as the models trained using real images when tested on real test datasets, they exhibit other advantages such as strong generalization capability. Concurrently, He et al. [8] also proposed to use large synthetic datasets for improving pretrained models on zero-shot and few-shot learning, and transfer learning.
@@
In this work, we propose to study the use of widely available pretrained generative models such as Stable Diffusion [25] as a source of additional labeled samples combinable with general class-incremental learning methods applied on large scale datasets.
@@ 3 Proposed method
The class-incremental training procedure is divided into steps: the base step, often referred to as step or initial step, followed by incremental steps. Each step consists of a training dataset containing images belonging to the set of new and previously unseen classes ; for . During each step, the model is trained on where is the replay memory containing few samples from previously encountered classes. After each step, the model is evaluated on the test set of all the classes learned so far without having access to any step or task descriptor. Additionally, during every step, the model also has access to an external source of data that may or may not change during the training. This external source of additional data can either be an online, and potentially infinite stream of data that the model can fetch when needed or a fixed set of data accessible in an offline manner.
@@ 3.1 Synthetic images
Following the recent in-depth study of Sariyildiz et al. [27], we propose to use a pretrained Stable Diffusion [25] as the source of additional data for incremental learning.
Prompting: To generate a sample for , the generative model Stable Diffusion requires a textual prompt. As proposed in [27], to generate a synthetic image for the class c we use the prompt “” where “” is the textual name of the class and “” its description. Although using only “c” as a prompt may be enough in most cases, this may lead to semantic errors with homographs. For example, in the CIFAR100 dataset using only “apple” as a prompt may generate images of products or shops from the homonym brand instead of the fruit. Adding the definition of the class to the prompt helps solve this issue. In order to produce these prompts automatically without any further engineering, WordNet [22] is used to provide the lemmas of the synset as the class name “c” and the definition of the synset as the description “”. For datasets such as ImageNet, the association of each class with a synset of WordNet is included in the dataset itself. For other datasets such as CIFAR100, the association with WordNet can be done semi-automatically at a negligible cost. Figure 1 shows synthetic samples for CIFAR100 before being resized to 32x32. In the cases where the mapping between the target dataset and WordNet can not be done, it may be necessary to either manually design and tune the prompt for each class or to rely on a different automation approach such as using image-to-text model for example.
Guidance scale: The classifier-free guidance [9] controls the trade-off between the quality of the generated synthetic samples and their diversity in diffusion models. Following Sariyildiz et al. [27], we also used a guidance scale of compared to the default to increase the diversity of our source of additional data at the cost of a slight decrease of the quality of each sample.
@@ 3.2 SDDR for Class-incremental learning
| iCaRL [24] | 57.17 | 52.57 | - | 65.04 | 59.53 | - | 51.36 | 46.72 |
| BiC [34] | 59.36 | 54.20 | 50.00 | 70.07 | 64.96 | 57.73 | 62.65 | 58.72 |
| LUCIR [10] | 63.42 | 60.18 | - | 70.47 | 68.09 | - | 64.34 | 61.28 |
| LUCIR w/ Mnemonics [20] | 63.34 | 62.28 | 60.96 | 72.58 | 71.37 | 69.74 | 64.54 | 63.01 |
| LUCIR w/ DDE [11] | 65.27 | 62.36 | - | 72.34 | 70.20 | - | 67.51 | 65.77 |
| LUCIR w/ BalancedS-CE [14] | 64.83 | 62.36 | 58.38 | 71.18 | 70.66 | 65.10 | 67.81 | 66.47 |
| TPCIL [30] | 65.34 | 63.58 | - | 76.27 | 74.81 | - | 64.89 | 62.88 |
| DER (w/o Pruning) [35] | 68.52 | 67.09 | - | - | 78.20 | - | - | - |
@@
Using the previously described generative model, we propose a new method for class-incremental learning: Stable Diffusion for Distillation and Replay (SDDR), which jointly uses the additional synthetic data with the real images from the replay memory .
@@
Datasets of real images used by concurrent approaches as an external source of additional data for class-incremental learning do not contain any image belonging to the same classes as the one encountered by the incremental model. This leads these methods to mainly use this external source of data in an unsupervised manner for knowledge distillation only. By leveraging the pretrained text-to-image generative model, our method is able to generate an additional dataset containing labeled images belonging to the same classes as the one previously learned by the model without the need for costly manual labeling and curating process. This allows us to use the additional dataset for distillation but also for replay in the classification loss. The ablation study in Section 4.3 shows the advantages of our approach.
@@
Our approach is designed as a complementary method that can be seamlessly combined with other standard methods for class-incremental learning. Algorithm 1 describes the training procedure of our approach SDDR when combined with a class-incremental learning model. Each step of the class-incremental learning procedure is divided into two distinct phases. During the first phase, the model is trained following the method it is combined with, using a mixed batch of data. As illustrated in Figure 2, each batch used to train the model is composed of half of synthetic images sampled from and the other half of real data sampled from the memory and the current dataset . Then in the second phase, before moving to the next incremental step, the synthetic dataset is extended by using Stable Diffusion to generate synthetic samples for each new class encountered in the current dataset , following the method detailed in Section 3.1. Therefore, at the end of each incremental step, the additional dataset would contain samples where is the number of classes encountered up to the step , included.
@@
In our work, we considered this additional dataset of synthetic samples as an offline dataset stored on the device itself. Nonetheless, if the storage is a constraint for the incremental learner, the additional dataset can be used in an online way as an infinite stream of data, similarly to [19], by either using the generative model locally or by relying on a cloud-based model. Compared to concurrent works relying on real datasets as a source of additional data, our method offers more flexibility depending on the constraint of the problem: it can be used with or without limited storage, computational budget, and capacity to communicate with external services.
@@
The quality of the synthetic images, their fidelity to the target classes, and the variety of concepts available in the additional external dataset are intrinsically limited by the pretrained generative text-to-image model used. Using a different and more specialized generative model or finetuning it on the specific target training dataset may further improve the performance of the incremental learning model it is combined with.
A naive and straightforward application of a pretrained text-to-image generative model in the context of class-incremental learning would be to use the synthetic datasets as a complete replacement of the replay memory for exemplar-free class-incremental learning. However, as discussed in the ablation study in Section 4.3, this approach does not achieve competitive results.
@@ 4 Experiments
| Methods | CIFAR100 | ImageNet-Subset | ||||
| =5 | 10 | 25 | 5 | 10 | 25 | |
| FOSTER [32] | 70.62 | 68.43 | 63.83 | 80.21 | 77.63 | 69.27 |
[…]
| FOSTER* [32] | 71.17 | 68.89 | 65.07 | 76.09 | 75.05 | 70.96 |
[…]
4.1 Experimental setups
Datasets: Experiments are conducted on three datasets: CIFAR100 [15], ImageNet-Subset and ImageNet [26]. CIFAR100 is composed of 60,000 32x32 RGB images from 100 classes with 500 training and 100 testing samples per class. ImageNet (ILSVRC 2012) is composed of around 1.28 million high-resolution images from 1,000 classes with about 1,300 training and 50 testing sample per class. ImageNet-Subset is a subset of ImageNet containing only 100 classes. Experiments on an alternative version of ImageNet-Subset can be found in the Supplementary materials. Following the experimental setting initially proposed by Hou et al. [10], an initial base step containing half of the classes is followed by (5, 10, or 25) incremental steps containing the remaining classes evenly divided. The class order is defined by NumPy using the random seed 1993. The training images are normalized, randomly horizontally flipped, and cropped, and no more augmentation is applied. For FOSTER [32], following its authors, AutoAugment [4] is also used for augmentation.
Comparison methods: We combine our proposed method SDDR with the recent state-of-the-art approach FOSTER [32] and with two fundamental baselines for class-incremental learning: iCaRL [24] and LUCIR [10]. We selected these two methods because they are representative of the principal approaches for class-incremental learning: using different distillation losses (logits distillation and feature distillation) and difference bias mitigation methods (Nearest-Mean Exemplar classifier and cosine normalized classifier). For comparison, we also include the recent approach PlaceboCIL [19] which leverages external unlabeled datasets of real high-resolution images to improve the distillation loss of class-incremental learning methods.
Implementation details: Following the standard setting [10, 24], a 32-layer ResNet [7] is used for CIFAR100 and a 18-layer ResNet for ImageNet and ImageNet-Subset. Every considered method uses a growing memory containing 20 exemplars per class during the training procedure. When combining SDDR with an existing baseline for class-incremental learning, the same hyperparameters as the original method are used. The only difference is that the batch size during incremental steps is effectively doubled by appending synthetic images generated by the generative model to the original batch of real images from the union of the new dataset and the replay memory. Some methods for class-incremental learning use other losses in addition to the classification and distillation losses. Using synthetic data for these other losses may improve or degrade the performance of the method on a case-by-case basis. Specifically, in our experiments on LUCIR [10], we decide not to use the synthetic data for the margin ranking loss as preliminary experiments suggested that it slightly decreases the performance of the method. Likewise, in our experiments on FOSTER [32], the synthetic data are used for the feature enhancement loss and the distillation loss during feature boosting and for the distillation loss during feature compression. We do not use the synthetic data for the logits aligned classification loss as it results in a significant decrease of the accuracy due to a bias of the classifier. For fair comparison with the other baselines, we report the accuracy of FOSTER after feature compression. For comparison of dual branch FOSTER before feature compression, see Supplementary materials. The generative model we use is Stable Diffusion version 1.4 [25] with a guidance scale of 2.0, 50 steps and the prompt “” [27]. Synthetic images have been generated using fixed seeds before the incremental training and every method is using the same datasets of synthetic images. The same seeds have been used for generating all classes, resulting in closely related classes having some similar images (same background, subject’s pose and position) where only the subject of the image changes. For CIFAR100, the synthetic dataset contains 500 32x32 images per class, and for ImageNet and ImageNet-Subset 1300 512x512 images per class. For every dataset, synthetic images are generated with size 512x512 and are only resized to 32x32 for CIFAR100 prior to the experiments.
Performance Measure: Models are evaluated and compared using the Average Incremental Accuracy defined by Rebuffi et al. [24]. It is the average of the Top-1 accuracy of the model on the test data of all the classes seen so far, at the end of each training step including the initial base step.
4.2 Comparison with baselines
@@
| LUCIR* [10] | 53.14 | 52.79 | 44.67 | 60.77 | 57.55 | 50.77 | 63.37 | 60.88 | 57.07 | 65.63 | 62.39 | 60.80 |
| LUCIR w/ PlaceboCIL [19] | 62.74 | - | - | 64.79 | - | - | 65.28 | - | - | - | - | - |
[…]
@@
In Table 1, the average incremental accuracy (Top-1) of our method and the different baselines are reported for CIFAR100, ImageNet-Subset, and ImageNet with 5, 10, and 25 incremental steps using a replay memory of 20 real exemplars per class for every method. Due to significant differences in the reproduced accuracy for FOSTER [32] between our work and Liu et al. [19], we decided to make a more detailed comparison with PlaceboCIL in Table 2.
@@
For FOSTER, LUCIR, and iCaRL, using our proposed approach SDDR leads to a significant improvement of the average incremental accuracy: from 1.01 percentage points (p.p.) to 4.54p.p. for FOSTER, from 0.87p.p. to 6.34p.p. for LUCIR and from 3.3p.p. to 8.44p.p. for iCaRL, depending on the dataset and the setting. This improvement is especially important in more challenging settings with numerous incremental steps or when the replay memory is small as discussed in Section 4.3. When looking more closely at the learned model, we can see that our approach improves the capacity of the model to preserve past knowledge without penalizing its plasticity. This is especially true in the most challenging settings with a small replay memory: for example on CIFAR100 with 10 incremental steps and 5 exemplars per class in the memory, by combining LUCIR with SDDR, the final accuracy on the 50 bases classes increases from 34.15% to 42.06% while the accuracy on the 50 remaining classes also increases from 46.08% to 52.23%. By combining our proposed method SDDR with FOSTER, we achieve new state-of-the-art performance on several datasets and settings.
Our proposed method achieves competitive performances compared to other considered approaches Mnemonics [20], DDE [11] or BalancedS-CE [14] and could be even combined with them to further improve their performance. Compared to PlaceboCIL [19], our method achieves similar or higher performance while using a significantly smaller dataset of additional external data: 50,000 synthetic images compared to the 1.28 million of real images from ImageNet for the experiments on CIFAR100 and 130,000 synthetic images compared to about 1.1 million of real images sampled without overlapping from ImageNet for the experiments on ImageNet-Subset.